OpenStudy (anonymous):

What is the surface area of this prism? A. 6480 m2 B. 3312 m2 C. 3240 m2 D. 1656 m2

OpenStudy (sleepyjess):

Just a sec

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

Ok, we need to do the same thing here, find the area of each individual shape, then add them up

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

For a triangle, you will use the formula \(\dfrac{h*b}{2}\) then multiply that by 2 since there are 2 triangles. The height for these triangles is 12 and the base is 18.

OpenStudy (anonymous):

so 12 x 18 x 2?

OpenStudy (anonymous):

divided by 2

OpenStudy (sleepyjess):

Actually, just 12*18 since you are dividing by 2 then multiplying by 2.

OpenStudy (sleepyjess):

It will give you the same result either way

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 216!

OpenStudy (sleepyjess):

Ok, now you need to find the areas of the rectangles.

OpenStudy (sleepyjess):

That would be 30*18 and (30*15)*2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 1980!

OpenStudy (sleepyjess):

Can you show me what you did so I can see what you did wrong?

OpenStudy (anonymous):

wait was i supposed to add them together...

OpenStudy (anonymous):

or leave them as individual answers...

OpenStudy (sleepyjess):

Yes, add them all together

OpenStudy (anonymous):

including the 216

OpenStudy (sleepyjess):

yep

OpenStudy (anonymous):

then i got 2196!

OpenStudy (sleepyjess):

Can you show me what you multiplied?

OpenStudy (anonymous):

was i wrong again... :/

OpenStudy (sleepyjess):

I may have miscalculated something, I just want to see where one of us went wrong.

OpenStudy (anonymous):

ok i multiplied 30*18 and 30*15*2 then i added the answers up with 216

OpenStudy (sleepyjess):

What did you get for 30*18?

OpenStudy (anonymous):

540

OpenStudy (sleepyjess):

Here is what I am showing \(30*18=540\\30*15=450*2=900\\540+900+216\)

OpenStudy (anonymous):

1656

OpenStudy (sleepyjess):

Yep

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

can you help me with like 3 more plz

OpenStudy (sleepyjess):

Yes

OpenStudy (anonymous):

ok thx